### v1

# LLM-Enhanced Static Analysis for Precise Identification of Vulnerable OSS Versions

Abstract.
Open-source software (OSS) has experienced a surge in popularity, attributed to its collaborative development model and cost-effective nature. However, the adoption of specific software versions in development projects may introduce security risks when these versions bring along vulnerabilities. Current methods of identifying vulnerable versions typically analyze and trace the code involved in vulnerability patches using static analysis with pre-defined rules. They then use syntactic-level code clone detection to identify the vulnerable versions. These methods are hindered by imprecisions due to (1) the inclusion of vulnerability-irrelevant code in the analysis and (2) the inadequacy of syntactic-level code clone detection. This paper presents Vercation, an approach designed to identify vulnerable versions of OSS written in C/C++. Vercation combines program slicing with a Large Language Model (LLM) to identify vulnerability-relevant code from vulnerability patches. It then backtraces historical commits to gather previous modifications of identified vulnerability-relevant code. We propose semantic-level code clone detection to compare the differences between pre-modification and post-modification code, thereby locating the vulnerability-introducing commit (vic) and enabling to identify the vulnerable versions between the patch commit and the vic. We curate a dataset linking 74 OSS vulnerabilities and 1013 versions to evaluate Vercation. On this dataset, our approach achieves the F1 score of 92.4%, outperforming current state-of-the-art methods. More importantly, Vercation detected 134 incorrect vulnerable OSS versions in NVD reports.
Open-source software security, Vulnerable version, Large Language Model
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Security and privacy Software security engineering
1. Introduction
Open-source software (OSS) has become increasingly popular in recent years, thanks to its collaboration and cost-effectiveness. In the rapidly evolving world of OSS, numerous versions exist due to continuous evolution. While OSS plays a pivotal role in expediting software development, the integration of particular software versions in development projects can pose security risks, as these versions may contain vulnerabilities. Therefore, having a comprehensive knowledge of vulnerable versions of OSS becomes imperative for software developers.
Public vulnerability repositories collect vulnerability reports of software products and disseminate information regarding the affected versions of the software. The National Vulnerability Database (NVD) (NVD, 2009), recognized as the largest public vulnerability database, employs the Common Platform Enumeration (CPE) format to store information about vulnerable versions. However, the NVD often encompasses all versions before the reported vulnerability or designates only the versions mentioned in the report as vulnerable. For instance, CVE-2018-5785 reported only version v2.3.0 as vulnerable in CPE (CVE, 2018). Yet, upon manual validation, it was uncovered that versions v2.1.1 to v2.3.0 are all susceptible to the vulnerability. Recent research (Dong et al., 2019; Mu et al., 2018; Tan et al., 2021; Nguyen and Massacci, 2013) indicates that incomplete and incorrect information about vulnerable versions is prevalent in such reports. Therefore, there is a need for an automated method that can more accurately identify vulnerable versions of released OSS vulnerabilities.
Two limitations of current approaches. One approach for confirming vulnerabilities involves utilizing Proof of Concept (PoC) to trigger the vulnerability. PoC triggers vulnerabilities through dynamic execution, offering conclusive evidence of their existence. However, executing a PoC for each software version is a time-intensive process that requires meticulous environment setup. Additionally, the PoC input may not be universal across all vulnerable versions (Dai et al., 2021). Therefore, current research has proposed the use of static analysis to identify vulnerable versions (Da Costa et al., 2016; Rosa et al., 2021; Nguyen et al., 2016; Rodríguez-Pérez et al., 2020; Xiao et al., 2020; Shi et al., 2022; Jang et al., 2012; Woo et al., 2022). These methods typically consist of two steps: extracting vulnerability features from code snippets and identifying vulnerable versions through code clone detection of these vulnerability features. The effectiveness of vulnerability feature extraction heavily relies on human expertise in formulating vulnerability rules. Existing methods typically extract vulnerability features using pre-defined patterns: some consider data and control dependencies of the patched code (Xiao et al., 2020; Shi et al., 2022); some consider using hashes of entire functions (Woo et al., 2021); and some consider only the patch code (Jang et al., 2012; Bao et al., 2022). These methods generally produce many false positives or false negatives because the features do not adequately characterize the vulnerability (❶). Existing methods for vulnerable code clone detection are limited to syntactic signatures (such as Levenshtein algorithm (Bao et al., 2022) and hash value comparison (Xiao et al., 2020; Woo et al., 2021)), without taking into account code changes that still preserve the semantics of functions (❷).
These limitations primarily stem from the lack of reasoning about the semantic context of the vulnerable code being analyzed. Recently, large language models (LLMs) have shown remarkable performance in code-related tasks including code generation (Du et al., 2024; Li et al., 2023) and code summarization (Ahmed et al., 2024; Ahmed and Devanbu, 2022). As such, we aim to address the first limitation (❶) by leveraging the power of LLM and in combination with static analysis. However, there is a challenge of Prompt engineering in leveraging a LLM in our problem domain. The design of prompts is pivotal in directing the LLM to produce desired responses. In vulnerability comprehension tasks, well-crafted prompts should encompass the code for analysis and offer a lucid description of the analysis objectives. However, in the case of current LLMs, merely providing all the vulnerable function codes and directly instructing vulnerability analysis often yields suboptimal outcomes. When the function code is too long, the LLM’s ability to understand the relationships between distant parts of the context may diminish (Li et al., 2024; Song et al., 2024). Hence, formulating prompts to guide LLMs in generating desired vulnerability logic analyses poses a challenge.
Due to variations in programming styles, variable naming, and structural differences, code clones may manifest different changes. Furthermore, developers often streamline and refactor code through function outlining. Such syntactic changes often mislead syntactic code clone detection-based approaches to incorrectly identify the vulnerable versions. This motivates us to propose a semantic-level code clone detection approach to address the second limitation (❷).
Our approach. We propose Vercation [^1] Vercation: Vulnerable version identification, a novel method to identify vulnerable versions of open-source C/C++ software by means of a symbiotic combination of static analysis, LLM, and semantic-level code clone detection. Given a vulnerability patch commit, Vercation applies program slicing to extract vulnerability-related statements as vulnerability features, leverages the capability of LLM in code understanding to refine the extracted features, and performs semantic-level clone detection on vulnerability features in function code changes. This hybrid method effectively overcomes the practical limitations of existing approaches. More specifically, Vercation automatically preprocess the patch commits and construct prompts based on the Few-shot and Chain-of-Thought (CoT) strategies, enabling the LLM to reason with the vulnerability and identify most probable vulnerable statements as features. Subsequently, Vercation traces earlier modifications of vulnerability features and applies a semantic clone detection method based on expanded Abstract Syntex Trees (ASTs) to pinpoint the vulnerability-introducing commit vic. Finally, Vercation identifies vulnerable versions between the vic and patch commit.
Evaluation. The first author meticulously curated a ground-truth dataset for evaluation, encompassing 11 commonly used Open Source Software (OSS) projects, 74 Common Vulnerabilities and Exposures (CVEs), and a total of 1,013 OSS versions. This dataset comprised every patch released through Git with respect to those 74 CVEs. The author engaged in manual examinations, meticulously labeling the presence of vulnerabilities across an extensive set of 10,476 software-version pairs. On this dataset, Vercation demonstrated both higher precision (89.2%) and recall (95.8%) compared to state-of-the-art methods (SOTAs), including V-SZZ (Bao et al., 2022) and V0Finder (Woo et al., 2021). We conducted an ablation study using two different LLMs (GPT-4 and CodeLlama) to evaluate their vulnerability comprehension capabilities. Utilizing the Few-shot and CoT combined strategy, GPT-4 achieved an F1 score of 92.4%, significantly outperforming the F1 score of Joern parser (Joe, 2014) by 44%, a commonly-used static analysis tool. More importantly, during the evaluation, we found 134 version errors in the NVD reports. Vercation has also shown to be efficient, analyzing each vulnerability on an average of 28.61 seconds.
Contributions. The main contributions of this paper are as follows:
  • We present Vercation, a novel, automated vulnerable version identification approach powered by LLM. Previous efforts heavily relied on pre-defined patterns of static analysis tools, while Vercation leverages the capability of LLM for vulnerability comprehension task, through the use of a multi-strategy universal prompt engineering.
  • Vercation presents a solution based on expanded AST to address the semantic-level clone detection challenge, which was not considered in existing techniques.
  • We meticulously assembled a comprehensive dataset that included 10,476 unique vulnerability-version pairs derived from 1,013 versions encompassing 74 CVEs. This extensive dataset was curated across 11 OSS projects and underwent meticulous labeling achieved through a combination of PoC input validation and rigorous manual verification.
  • We have implemented a prototype of our approach and assessed its performance using our dataset, achieving the F1 score of 92.4%, outperforming SOTAs by 9% to 48%. More importantly, by applying our approach, we have detected 134 incorrect vulnerable OSS versions in NVD reports.
The source code of Vercation and our curated dataset is publicly available at https://anonymous.4open.science/r/Vercation_A.
2. Background and Motivation
In this section, we clarify the target problem of Vercation, introduce the LLM and discuss the motivation with two examples.
2.1. Problem Statement
Commits serve as comprehensive records of OSS development, functioning as vital checkpoints for tracking the chronological evolution of code changes. They enable developers to revisit specific points on a particular date or time. This paper centers on vulnerable version identification in OSS vulnerabilities. Within the development of an OSS, certain commits introduce vulnerabilities, referred to as vulnerability-introducing commits (vic), which are later fixed in patch commits (pc). The vulnerable version analysis aims to pinpoint the vic, allowing us to assess which versions are susceptible to the vulnerabilities.
We posit that the initial function undergoes modification by a vic and transforms into a vulnerable function $F_{v}$. There probably exists some subsequent commits such as feature-adding commits and refactoring commits to optimize the function code, which is denoted as $F_{vr}$. Ultimately, software analysts discover the vulnerability and patch the code in the pc, resulting in the final function $F_{p}$. An illustrative timeline is presented in Figure 1.
Figure 1. Function Timeline from Vulnerability Introduction to Vulnerability Fixing.
2.2. Large Language Model
Large Language Models (LLMs) have revolutionized natural language processing and have demonstrated remarkable capabilities in various code-related tasks. These models, extensively trained on vast amounts of text and code repositories, can understand and generate human-like text and code. Recent advancements in LLMs, such as GPT-3 (Brown et al., 2020), GPT-4 (Achiam et al., 2023), and CodeLlama (Roziere et al., 2023), have demonstrated impressive performance in code generation (Du et al., 2024; Li et al., 2023), code summarization (Ahmed et al., 2024; Ahmed and Devanbu, 2022) and program repair (Fan et al., 2023; Hossain et al., 2024)..
The success of LLMs in code-related tasks can be attributed to their ability to capture complex patterns and semantic relationships in code. This makes them particularly suitable for tasks that require deep understanding of code structure and functionality like vulnerability detection. However, the effectiveness of LLMs heavily depends on the quality of prompts used to guide their responses. Prompt engineering, the process of designing effective prompts, has become a crucial area of research in leveraging LLMs for specific tasks (Zhou et al., 2022; Chen et al., 2023).
Despite their power, LLMs also face challenges in code analysis tasks. These include potential biases in their training data and difficulties in handling very long code sequences. In our work, we leverage the semantic understanding capabilities of LLMs to enhance the accuracy of static analysis tool in vulnerability-related code extraction, a novel application in the field of software security analysis.
2.3. Motivating Examples
We present two examples of disclosed vulnerable code in two distinct scenarios, which motivate our work. Both of these vulnerabilities have now been effectively resolved by the development team. The key points we wish to underscore are as follows: 1) The significance of understanding vulnerability behavior in the process of discovering vic. 2) The challenge of accurately pinpointing vic due to the significant syntax changes introduced by a refactoring commit.
1diff --git a/libavformat/mxfdec.c b/libavformat/mxfdec.c
2index 70142c4..012d8f5 100644
3--- a/libavformat/mxfdec.c
4$+$ b/libavformat/mxfdec.c
5@@ -493,11 +493,11 @@ static int mxf\_read\_primer\_pack
6static int mxf_read_primer_pack(void *arg, AVIOContext *pb, int tag, int size, UID uid, int64_t klv_offset)
7{
8 MXFContext *mxf = arg;
9 int item_num = avio_rb32(pb);
10 int item_len = avio_rb32(pb);
11 if (item_len != 18) {
12 avpriv_request_sample(pb, "Primer pack item length %d", item_len);
13 return AVERROR_PATCHWELCOME;
14 }
15-  if (item_num > 65536) {
16+  if (item_num > 65536 || item_num < 0) {
[…]
18 return AVERROR_INVALIDDATA;
19 }
20 if (mxf->local_tags)
[…]
22 av_free(mxf->local_tags);
23 mxf->local_tags_count = 0;
24 mxf->local_tags = av_calloc(item_num, item_len);
25 if (!mxf->local_tags)
26 return AVERROR(ENOMEM);
27 mxf->local_tags_count = item_num;
28 avio_read(pb, mxf->local_tags, item_num*item_len);
29 return 0;
Listing 1: Motivating Example of CVE-2017-14169.
Example 1) A vulnerability in FFmpeg (CVE-2017-14169 (CVE, 2017b)) shows the limitation of existing approaches in terms of capturing vulnerability logic. The code snippet of pc is shown in Listing LABEL:list:CVE-2017-14169. A sanitizing check was added at Line 16 for the item_num. Without this check, a remote attacker can make a crafted file with a large item_num field such as 0xffffffff, causing a buffer overflow issue in the avio_read function at Line 28 and potentially leads the application to exhibit incorrect behavior or crash.
V-SZZ (Bao et al., 2022) assumes the deletion lines in the security patch as vulnerable codes and the basic idea of it is to pinpoint the earliest commit introducing the deletion lines as vic. However, vulnerability logic is composed of various vulnerable codes, such as Lines 24 and 28 in Listing LABEL:list:CVE-2017-14169. V-SZZ only considers tracing back deleted line (Line 15) and overlooks the real vulnerability behavior, leading to incorrect identification of vic. V0Finder (Woo et al., 2021) attempts to identify vulnerable versions using code clone detection. It generates hash values for the entire vulnerability function and identifies code clones according to the distance value of two hashes. This method may still introduce excessive vulnerability-unrelated features due to the inclusion of “entire function”.
Observation. Identification of vulnerable versions is hindered by the challenge of comprehending vulnerability logic from the security patch. LLMs can grasp contextual code semantics and achieve human-like understanding in code-related tasks, eliminating the reliance on predefined patterns and rules utilized by traditional static analysis and automated tools. Through the integration with an LLM, we can automate vulnerability analysis, enhancing the efficiency of our approach.
1diff --git a/libavformat/avidec.c b/libavformat/avidec.c
2index 57fbcd7a1b341..bab62a08c1c1d 100644
3--- a/libavformat/avidec.c
4$+$ b/libavformat/avidec.c
5@@ -350,8 +350,7 @@ static void avi_read_nikon(AVFormatContext *s, uint64_t end)
6 uint16_t tag = avio_rl16(s->pb);
7 uint16_t size = avio_rl16(s->pb);
8 const char *name = NULL;
9 char buffer[64] = { 0 };
10- if (avio_tell(s->pb) + size > tag_end)
11-    size = tag_end - avio_tell(s->pb);
12+ size = FFMIN(size, tag_end - avio_tell
13+    (s->pb));
14 size -= avio_read(s->pb, buffer, FFMIN(size,
15 sizeof(buffer) - 1));
Listing 2: Motivating Example of Semantic-Invariant Modification in Refactoring Commit.
Example 2) After extracting the vulnerability logic and its corresponding vulnerable statements ($S_{v}$), we can locate the vic by backtracing the code changes of $S_{v}$ in the previous commit, thereby identifying the vulnerable versions. V-SZZ (Bao et al., 2022) identifies code clones before and after such code changes by the edit distance. However, refactoring commits, prevalent in OSS development, aims to improve and reorganize code while preserving its functionality. Such commits can result in situations where the code’s semantics remain the same, but there is a significant syntactic difference between the code before and after modifications. Essentially, such a commit would cause the tool like V-SZZ to incorrectly flag it out as a vic.
To exemplify such a scenario, we present code changes made in the file avidec.c of FFmpeg. As depicted in Listing LABEL:semantic, the added lines convey identical semantic meaning with deleted lines, which both represent selecting the minimum value between size and tag_end-avio_tell(s->pb) and assigning it to the variable size.
Due to the refactoring of the if structure into the FFMIN function, the edit distance utilized by V-SZZ (Bao et al., 2022) perceives a significant syntactic difference between the two (48% similarity), leading to incorrect pinpoint of the vic. ReDeBug (Jang et al., 2012) and V0Finder (Woo et al., 2021) relies on syntax-based pattern matching for code clone detection. MVP (Xiao et al., 2020) and AFV (Shi et al., 2022) extract the dependencies between two statements and normalize them to hash values as function signatures, overlooking the semantic information of statements.
Observation. This case emphasizes the significance of identifying semantic invariance in the code modifications, which was often overlooked in existing techniques. AST is a tree representation of code that explicit statement order and the execution logic, which contains rich semantic information (Yang et al., 2021). Vercation introduces a code clone detection grounded in expanded AST analysis.
Figure 2. Vercation Workflow.
3. Design of Vercation
We propose Vercation, an end-to-end automated approach designed for vulnerable version identification in OSS vulnerabilities. The high-level workflow is illustrated in Figure 2. Vercation consists of three phases: Vulnerable code extraction (P1), Code changes detection (P2) and Vulnerable version range delineation (P3).
In P1, Vercation combines program slicing and LLM to identify and extract vulnerability-related program statements in a precise manner. Specifically, we utilize the patch code as a slicing criterion to extract dangerous flows. To mitigate the risk of including statements unrelated to the vulnerability (false positives), we employ prompt construction with few-shot and Chain-of-Thought (CoT) strategies, empowering LLM to reason with the vulnerability based on the extracted dangerous flows and accurately extract vulnerability-related statements.
In P2, Vercation retraces historical commits to collect previous modifications of vulnerable statements. For each statement before and after modification, Vercation expands the functions within the statement, generates ASTs and normalizes them. Then we utilize an in-order traversal algorithm to compare the AST before and after the modification as semantic similarity, which determines if the commit is the initial introduction of the vulnerable statements.
In P3, Vercation identifies the versions affected by a vulnerability based on the CVE’s pc and the corresponding vic.
In the following sections, we explain these phases in detail.
3.1. Vulnerable Code Extraction
Taking a vulnerability patch commit as an input, Vercation identifies and extracts program statements related to vulnerability in three steps as explained in the next three subsections respectively.
3.1.1. Dangerous flow extraction
Firstly, Vercation conducts program slicing (Weiser, 1984) on the source code to extract crucial program statements that may contribute to the vulnerability, which we call them as dangerous flow. The slicing is based on the following criteria: deleted/added statements and patch-related variables. Building upon existing techniques (Xiao et al., 2020; Shi et al., 2022; Woo et al., 2022), we execute program slicing on the program dependency graph of the patched function according to the slicing criteria. For different statement types, we customize different slicing rules:
  • Assignment statement affects the data-flow values of vulnerability behavior. We conduct normal slicing in the assignment statements and the assigned variables should be added to the slicing criterion. For example, if we take variable item_num of Line 15 and 16 in Listing LABEL:list:CVE-2017-14169 as slicing criterion and perform forward slicing, Line 24 and 27 are included.
  • Conditional statement affects the reaching condition of vulnerability trigger. We aim to slice the condition statement that takes the variables of slicing criterion as condition check, and the result also includes all the subsequent statements of condition statement (e.g., subsequent statement Line 17-18 of condition statement Line 16, subsequent statement Line 26 of condition statement Line 25).
  • Function call statement. If the function call statement’s parameters contain patch-related variables included in the slicing criterion, we conduct slicing on the statement.
  • Return statement. There is no need for forward slicing because there is no dependency between the return value and the statements following the return statement. For example, there is no need for forward slicing on Line 18 and 26 in Listing LABEL:list:CVE-2017-14169.
Meanwhile, the slicing is performed in two directions: backward and forward. Backward slicing is to trace the source of patch-related variables. For example, when we designate item_num (e.g., Line 15, 16 in Listing LABEL:list:CVE-2017-14169) as the criterion, the result of the backward slice is Line 9 (Line 15 is data-dependent on Line 9). Additionally, Line 15 is control-dependent on Line 11, as its execution is contingent on the unsatisfied condition in Line 11; otherwise, the function will promptly return. Following backward slicing, we retain Lines 9 and 11 as backward dangerous flow. Forward slicing is used to find the trigger behavior that causes the vulnerability. For example, in Listing LABEL:list:CVE-2017-14169, if we set the modified statements Lines 15 and 16 along with the variable item_num as criteria, the outcomes of forward slicing include 1) Lines 24, 27 and 28, which are directly or indirectly data-dependent on the variable item_num and 2) Lines 17, 18, 25 and 26, which are directly or indirectly control-dependent on the variable item_num.
3.1.2. Dangerous flow-based prompt construction
Secondly, we leverage an LLM to refine these dangerous flows further since the dangerous flows extracted in the above step (program slicing) may include false positives (unrelated to vulnerability logic). We explore prompting strategies to assist LLM in vulnerability comprehension task and in refining vulnerable statements from the dangerous flows extracted through program slicing.
Our method utilizes an LLM that supports interaction with system prompts and user prompts. The system prompt sets the role and background of the LLM. The user prompt consists of specific instructions issued by the user. At the beginning of the system prompt, we assign the role of a security researcher to the model with the statement, ”You are a security researcher, an expert in detecting security vulnerabilities,” and indicate that we will provide the CVE information and dangerous flow. The system prompt concludes with a declaration of the fixed structure expected for the model’s response. The content of the user prompt includes detailed CVE information including CVE ID, CWE ID, CVE description, and dangerous flow extracted from program slicing.
Additionally, we use the following strategies to determine the features of the user prompt:
1) Few-shot prompting. Few-shot prompting is a technique where we provide the LLM with a small number of examples demonstrating the desired task before presenting the actual problem. This approach helps guide the model’s behavior by showing it how to respond to similar inputs. We include two carefully crafted examples in the user prompt before providing the target CVE information. Each example consists of: a) A sample CVE ID, CWE ID and CVE description. b) Corresponding dangerous flow extracted from program slicing. c) The expected vulnerability logic analysis. d) The correctly identified vulnerable statements.
2) Chain-of-thought prompting. Chain-of-thought (CoT) prompting provides LLM with a prompt that encourages it to generate intermediate reasoning steps before arriving at a final answer (Wei et al., 2022). We prompt the model to ”Please reason the vulnerability logic from the provided code” before concluding the vulnerable statements.
Our prompting template is shown in Figure 3 and we take the vulnerable statements in the answer as the input of the next step.
System:You are a security researcher, expert in detecting security vulnerabilities.I will provide you with a CWE ID, CWE ID, CVE description, and dangerous code.Please extract the vulnerability logic from the code and indicate which statementsare relevant to the vulnerability logic.Provide a response only in the following format: vulnerability logic: ‹text¿vulnerable lines : [¡Line number List¿]Do not include the added line number (with +) and anything else in response.User:¡CVE ID¿,¡CWE ID¿,¡CVE description¿ and ‹dangerous flows¿
Figure 3. The LLM Prompt Template.
3.1.3. Statement weight allocation
As vulnerable statements contribute differently to vulnerability triggers, we propose strategies to allocate weights to the vulnerable statements extracted by LLM. We have noticed that certain types of C/C++ vulnerability often involve specific functions. For instance, in Listing 1, Lines 24 and 28 are critical trigger lines. With this knowledge, we can rely on a pre-defined table (Table 1) to identify such functions within vulnerable statements extracted by LLM. We summarized the common types of vulnerabilities in C/C++, and combined the authors’ domain knowledge with the enumeration of sensitive functions from relevant papers (Shu et al., 2017; Alaba et al., 2017) to create the pre-defined table.
Some OSS commonly used customized functions serve as variants of sensitive functions. For instance, the av_calloc function on Line 24 is a variant of the calloc function. To discern whether these customized functions contain sensitive functions, we conduct inter-procedure analysis. Taking Listing LABEL:list:CVE-2017-14169 as an example, considering this is an overflow-type vulnerability, we recognize the av_calloc in Line 24 and avio_read function in line 28 as sensitive functions.
Parameters. There are two parameters for the weight allocation of statements, i.e., weight_v and weight_d, that could affect the effectiveness of Vercation: 1) weight_d is the default weight for initializing the weights of all statements. 2) weight_v is the weight for the statements containing dangerous functions. We set weight_v=2$\times$×weight_d, weight_d=1.0. Note that these weights are used in Code changes detection phase (P2).
3.2. Code Clone Detection
Given the vulnerable statements $S_{v}$ of a patch commit (extracted in Section 3.1), Vercation performs a commit backtrace. The goal is to identify which previous commit introduced $S_{v}$ by tracing the history of earlier modifications. To achieve this, Vercation employs the git blame command to backtrack through previous commits. We introduce a semantic-level similarity comparison method to automatically detect code clones before and after modifications. We define the statements of the post-modification commit as $S$, and those from the pre-modification commit as $S^{\prime}$, with $S_{v}\subseteq S$.
We first use line mapping to pre-filter highly similar lines in $S^{\prime}$ with $S_{v}$. If such lines are found, it indicates that the commit is not the first to introduce $S_{v}$. We use the edit distance to calculate line similarity $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})$, $S_{i}\in S^{\prime}$. We set a similarity threshold $\vartheta_{1}$, if $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})\geq\vartheta_{1}$, we continue to backtrack the previous commit. If $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})<\vartheta_{1}$, we proceed to make a comparison with the next semantic matching.
Table 1. Sensitive Functions for Some Vulnerability Types.
Vulnerability Type Sensitive Function
Buffer overflow strcpy, strncpy, memcpy, memset, read, write, gets, gets_s, strcat
Integer overflow add, multiple, bit-shifting, memcpy
Null pointer derefrence malloc, calloc, realloc, strdup
Pointer out-of-bounds access memcpy, memmove
Use after free free, malloc, calloc
Format string sprintf, printf, scanf
Arbitrary command execution fopen, popen, system
3.2.1. AST-level semantic comparison
As exemplified in example 2 of Section 2.3, Lines 5-6 (deleted pre-commit lines) and 7-8 (added line of post-commit, e.g., included in $S_{v}$) exhibit high semantic similarity. However, previous research failed to identify this high similarity through simple syntactic analysis, resulting in the commit (version) being erroneously labeled as an introducing commit (vulnerable version). We propose a semantic similarity comparison method at AST-level to address this limitation. AST is a tree representation of code that preserves well-defined components of statements, explicit statement order and the execution logic, which contains rich semantic information (Yang et al., 2021).
1) Function inlining. Due to code refactoring often integrating code snippets into a new function, as shown in Listing LABEL:list:false_reason2, the codes in Lines 7-9 are integrated into functions SSL2_RECORD_add_length, SSL2_RECORD_add_off and SSL2_RECORD_get_length. Therefore, we utilize the inline technique to expand functions during AST generation to gather more semantic information about the statements. This involves obtaining the function definition at the callsite and expanding the function body inline, resulting in a more comprehensive code representation. We get the AST for the target file using Clang without performing actual compilation. Furthermore, through recursive retrieval of line numbers for sub-ASTs, we can locate the sub-AST corresponding to $S_{v}$.
1@@ -567,18 +569,18 @@ int dtls1_read_bytes(SSL *s, int type, unsigned char *buf, int len, int peek)
2%diff --git a/ssl/record/rec_layer_d1.c b/ssl/record/rec_layer_d1.c
3%index 80f10be480222..c58af7f36abee 100644
4%--- a/ssl/record/rec_layer_d1.c
5%$+$ b/ssl/record/rec_layer_d1.c
6 if (!peek) {
7- rr->length -= n;
8- rr->off += n;
9- if (rr->length == 0) {
10+ SSL3_RECORD_add_length(rr, -n);
11+ SSL3_RECORD_add_off(rr, n);
12+ if (SSL3_RECORD_get_length(rr) == 0) {
13 s->rlayer.rstate = SSL_ST_READ_HEADER;
14
15%In ssl/record/record_locl.h
16#define SSL3_RECORD_add_length(r, l) ((r)->length += (l))
17#define SSL3_RECORD_add_off(r, o) ((r)->off += (o))
18#define SSL3_RECORD_get_length(r) ((r)->length)
Listing 3: Example of Code Outlining Modification.
2) AST normalization. In programming, different coders may have different coding styles, resulting in variations in code structure. To ensure consistency and minimize noise introduced by coding style differences, it is imperative to normalize the AST. AST normalization involves conditional structure normalization, loop structure normalization and relational operation normalization.
  • Conditional structure can be implemented by if-else and switch-case. We normalize the structure of switch-case: switch(expression){case value1: statement1; break; case value2: statement2; break; default statement3;} into a if-else structure: if(exp == value1) statement1; else if(exp == value2) statement2; else statement3.
  • Loop structure can be represented as while, do-while and for structures, we normalize all the loop structure to while (Xue et al., 2022). The for structure is for(initialize; condition; increment) {statements;}, we transform it into while structure: initialize; while(condition) {statements; increment}.
  • Relational operation include commutative and noncommutative operation. Commutative relational operation (i.e., +, !=, &&) refers to the operation where commutate the operands does not change the operation semantics. We normalize the two operands (left and right child nodes) of the operation (parent nodes) by alphabetical order. For example, we transform $b+a$ (b is the left child node) to $a+b$ (a is the left child node). Noncommutative relational operation (i.e., $>$, <=<=< =) will change the operation semantic when commutating the operands, we normalize them by predefined rules. For example, we transform $b<a$ to $a>b$ (normalize operation $<$ to $>$).
1
Input : Post-modification statement’s $AST_{v}$, Pre-modification statements’s $AST_{i}$
Output : Boolean indicating if the commit introduces vulnerabilities
2
3Function ASTSimilarity($AST_{v},AST_{i}^{{}^{\prime}}$):
4       $Sq_{v},Sq_{i}=inOrderTraversal(AST_{v}),inOrderTraversal(AST_{i}^{{}^{\prime}})$
5       $Similarity=EditDistance(Sq_{v},Sq_{i}^{{}^{\prime}})$
6       if $Similarity\geq\theta$ then
7             $return\ True$
8      else
9            $return\ False$
10       end if
11      Function inOrderTraversal($AST$):
12             if $AST\ is\ null$ then
13                   $return\ empty\ list$
14             end if
15            $return\ inOrderTraversal(AST.left)\ +\ [AST.node]\ +\ inOrderTraversal(AST.right)$
16
17
18
Algorithm 1 AST-based Similarity for Code Changes
3) AST similarity. Ultimately, we utilize an in-order traversal algorithm to ascertain whether the commit introduces $S_{v}$ by comparing the AST similarity of $S_{v}$ and $S_{i}^{{}^{\prime}}$, as shown in Algorithm 1. First, we perform an in-order traversal of the ASTs for $S_{v}$ and $S_{i}$, converting the results into sequences $Sq_{v}$ and $Sq_{i}^{{}^{\prime}}$. Then, we calculate the edit distance between the two sequences as their semantic similarity $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})$. We set a similarity threshold $\vartheta_{2}$, and if $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})\geq\vartheta_{2}$, $S_{v}$ and $S_{i}^{{}^{\prime}}$ are considered to be semanticly similar. If $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})<\vartheta_{2}$, it indicates that there is no statement similar to $S_{v}$ in the pre-modification commit.
3.2.2. Code clone determination
We calculate the similarity_score for each code modification of the previous commit with the backtrace. We set a threshold $\vartheta_{3}$, if $similarity\_score<\vartheta_{3}$, we determine that there is no code clone between the pre-modification code of this previous commit and the vulnerable statements, thus confirming that this commit introduced the vulnerability. Otherwise, we proceed with the commit backtrace process.
$\begin{split}similarity\_score=\frac{sim\_a\times weight\_v+sim\_b\times weight\_d}{a\times weight\_v+b\times weight\_d}\end{split}$
where, $a$ and $b$ respectively refer to the counts of all the statements of $S_{v}$ in two different weight types. sim_statement refers to the statements where $Similarity_{A}(S_{v},S_{i}^{{}^{\prime}})\geq\vartheta_{1}$ or $Similarity_{B}(Sq_{v},Sq_{i}^{{}^{\prime}})\geq\vartheta_{2}$ (i.e., semantically or syntactically similar) during each commit comparison. $sim\_a$ and $sim\_b$ respectively refer to the sim_statement counts in two different weight types.
3.3. Vulnerable Version Range Delineation
Since vulnerabilities introduced by vic will be transferred to future commits until they are patched, we identify the versions affected by a vulnerability based on the CVE’s patch commits and the corresponding vic. During the development process of a GitHub project, developers assign tags for different versions. The vulnerable version tags are located between the vic and patch commit.
If commit A is reachable to another commit B, it indicates that A belongs to the parents or the ancestor commits of B (Bao et al., 2022). We first identify the version tags reachable to vic (i.e., $V_{i}$). Then, we identify the version tags reachable to pc (i.e., $V_{p}$). Finally, we obtain the vulnerable version set $V_{v}$, where $V_{v}=V_{i}-V_{p}$. For example of the CVE-2021-20294 (CVE, 2021a), the version tags reachable to pc are [binutils-2_36, binutils-2_42], and the version tags reachable to vic are [binutils-2_35, binutils-2_42], binutils-2_42 is the latest version tags of Binutils project. Therefore, we consider the version tags [binutils-2_35, binutils-2_36) as vulnerable.
4. Experimental Setup
4.1. Supporting Tools
We utilized Joern (Joe, 2014) for program slicing and developed Python scripts to extract control and data dependencies, outputting the sliced statements as dangerous flows. We explored two LLM models, GPT-4 and CodeLlama-13B, for vulnerable code extraction. We use the public API developed by OpenAI to perform the experiment in GPT-4. We set the temperature parameter to 0 to obtain conservative and deterministic answers and the maximum size of tokens to 1024. We deployed the Codellama model on our server with four NVIDIA RTX A5000 GPUs. Similar to GPT-4, the temperature parameter is set to 0, and the maximum size of tokens is set to 1024. Subsequently, we apply weight allocation strategies to extract vulnerable statements using Python scripts. In the code changes detection, we utilize the Clang 10.0.0 tool for function expansion (inline optimization) and AST generation of C/C++ code. Specifically, we use the -fsyntax-only option to perform syntax checking without actual compilation. We then develop Python scripts to achieve AST normalization and similarity comparison. Overall, we constructed our system with 6K LoC in Python.
4.2. Dataset
With confirmation from the author of V-SZZ, it is established that V-SZZ assumes the deleted lines in the patch commit as vulnerable codes and manually labels the first introduction of deleted lines in a patch commit as the vic to construct the dataset. However, as mentioned in Section 2.3, deleted lines cannot fully capture the vulnerability logic. Additionally, it cannot be guaranteed that the identified software version will necessarily exhibit the vulnerability. Therefore, we need to construct a ground-truth dataset with labeled vulnerable version or non-vulnerable version of OSS, which are verified by public PoC.
Vulnerabilities. Our dataset comprises 74 publicly disclosed CVEs associated with public PoC and patches, containing 11 different C/C++ OSS in Table 2. These CVEs span 11 prevalent OSS, as detailed in Table 2. The selected OSS is widely used and has a certain number of versions. Meanwhile, the collected vulnerabilities cover 13 common CWE types.
Patches. We acquired security patches by crawling patch commits from the OSS Git repositories. In total, the patches cover 1,357 modification lines containing 941 insertion lines and 416 deletion lines. Among these, 246 lines are unrelated to vulnerability (e.g., modifications in the ChangeLog file). As show in Table 2, on average, each patch commit has 21.22 modification lines. The number of added lines significantly exceeds that of deleted lines, underscoring the critical importance of incorporating additional code into the analysis process. Among these, 12.5% are large patch commits (the number of modification lines exceeding 50). Notably, 47 patches (63.51%) contain both code insertion and deletion, 21 patches exclusively insertions (28.38%), and 6 patches solely involve deletions (8.11%).
Verifying & Labeling. We apply the following process to systematically label the ground-truth dataset:
1) The fixed version can be determined through the official release of patch commits. We confirm whether versions following the fixed version are vulnerable by checking if they contain the patch code.
2) For the remaining version, we obtained the public PoC of each vulnerability from the Git issues or referring to OSS sites.
3) To validate the PoC on each corresponding OSS version, we establish the dependent environment for each OSS version and compile the libraries.
4) During the PoC input testing process for each vulnerability, we analyze the triggering conditions and dangerous behaviors of the vulnerability and label the statements that trigger the vulnerability as $S_{trigger}$. For versions where the PoC cannot trigger the vulnerability (as some PoCs are only applicable to certain specific versions), we check whether the version contains $S_{trigger}$ to label the version.
Table 2. Ground-truth Dataset Overview.
IDX Name #Version #CVE #Average Diff Lines Domain
1 Binutils 139 11 14.64 (+12.09/-2.55) Programming tools
2 cJSON 46 1 30.00 (+29.00/-1.00) JSON parser
3 FFmpeg 382 18 21.72 (+12.72/-9.00) Multimedia processing
4 Libarchive 42 9 16.83 (+14.00/-0.86) Streaming processing
5 Libcaca 7 3 36.67 (+25.67/-11.00) Graphics library
6 Liblouis 62 6 4.67 (+3.17/-1.50) Braille translator
7 Libming 14 9 12.22 (+9.78/-2.44) Flash library
8 Libtiff 75 6 16.17 (+13.50/-2.67) Image tools
9 Libxml2 204 3 43.00 (+18.75/-24.25) XML toolkit
10 OpenJPEG 22 5 17.80 (+13.40/-4.40) Image codec
11 Pcre2 20 3 21.67 (+15.00/-6.67) Regular expression engine
Total - 1013 74 21.22 (+15.19/-6.03) -
4.3. Baseline and Metrics
Baseline. For the vulnerable version identification task, we selected the NVD (CPE, 2009), along with SOTAs: SZZ algorithms (AG-SZZ (Kim et al., 2006), B-SZZ (Falleri et al., 2014), V-SZZ (Bao et al., 2022)) and V0Finder (Woo et al., 2021) for accuracy comparison. We gathered the vulnerable version range of each CVE from CPE to evaluate the NVD’s accuracy. The basic idea of SZZ algorithms is to backtrace the commit history to locate the earliest commit that introduces the deletion lines removed by security patches. We use different SZZ algorithms to identify the vic, with the versions between the vic and the pc considered vulnerable. Since these algorithms trace deleted lines individually, instances may arise where multiple vic are identified, and in such cases, we select the earliest one as the detection outcome. V0Finder is a vulnerability introduction detection tool that uses a clone-based technique. We first use V0Finder to generate vulnerable function fingerprints for our ground-truth dataset, and then generate fingerprints for all functions of all OSS versions. Subsequently, we use V0Finder’s clone-based approach to identify vulnerable versions. For Vercation, We selected thresholds $\vartheta_{1}=0.9$, $\vartheta_{2}=0.8$ and $\vartheta_{3}=0.7$ (related experiments are introduced in Section 5.3).
In terms of the performance of vulnerability code extraction, we compared the differences between combining various LLMs (GPT-4, CodeLlama-13B) and using the static analysis tool Joern parser (Joe, 2014) alone.
Metrics. To evaluate the accuracy, we employ the 3 following three metrics, i.e., true positives (TP), false positives (FP), false negatives (FN), precision (#$\frac{\#TP}{\#TP+\#FP}$), recall (#$\frac{\#TP}{\#TP+\#FN}$) and F1-score ($\frac{2*precision*recall}{precision+recall}$), to measure the accuracy of the above methodologies. These metrics are also used in previous studies to evaluate the technique performance (Bao et al., 2022; Rosa et al., 2021). Because of significant differences in the version count for distinct OSS, we compute the precision and recall for each OSS and report the average of precisions and recalls finally.
5. Evaluation
In this section, we answer the following research questions to evaluate the performance of the proposed Vercation.
RQ1. Is Vercation more effective than the state-of-the-art vulnerable version identification methods?
RQ2. How effective is Vercation powered by LLMs in vulnerability comprehension?
RQ3. How accurate is our proposed semantic-level code clone detection method?
RQ4. What are the effects of applying Vercation in the real world?
5.1. RQ1: Effectiveness of Vercation
Table 3 shows the vulnerable version identification results of NVD, SZZ algorithms, V0Finder and Vercation. Note that some CVEs do not provide vulnerable versions in NVD, such as CVE-2021-30499 (CVE, 2021b). Meanwhile, SZZ algorithms cannot work on the patch commits containing solely added lines (i.e., 11 cases), so we exclude these cases. In the Vulnerable Code Extraction phase of Vercation, we utilize the optimal LLM (GPT-4) along with the most effective prompt strategy (Few-shot+CoT), which are comprehensively detailed in Section 5.2. Vercation identified 1612 correct vulnerable versions with 89.2% precision and 95.8% recall, outperforming than the other SOTAs.
5.1.1. Comparison with NVD
NVD reports 224 true positives, 333 false positives and 1398 false negatives with 69.9% precision and 42.4% recall.
1) The reason for false positives in NVD is that it sometimes assumes that all previous versions of disclosed vulnerabilities are vulnerable, without conducting a specific analysis. This situation exists in 12 CVEs (16%) of the ground truth. For instance, security experts found a buffer overflow risk exists in Libtiff 4.4.0, but NVD simply reported the versions of Libtiff$<$4.4.0 were exposed in vulnerability. In reality, the vulnerable version range is Libtiff$<$4.4.0 and $\geq$4.0.0.
2) The reason for low recall in NVD can be summarized into two types: i) NVD only reported the version in which vulnerabilities were found. For example, CVE-2017-14152 (CVE, 2017a) was found in OpenJPEG 2.2.0, then NVD reported the affected version was only 2.2.0, whereas the real vulnerable version range was 2.2.1-2.2.0. The problem exists in 45 CVEs (61%) of the ground truth. ii) Some vulnerabilities are not fixed promptly after disclosure but are addressed after several versions have been released. These vulnerable versions can be overlooked if NVD does not update the version information. For example, CVE-2021-33815 (CVE, 2021c) was discovered in FFmpeg 4.4 but was fixed in version 5.0. Due to the lack of timely tracking of the fix information, NVD overlooked the vulnerable versions 4.4.1-4.4.4.
Table 3. Accuracy of Vulnerable Version Identification.
Methods Reference Correctvul Precison Recall F1 score
AG-SZZ  (Fan et al., 2019) 962 0.405 0.627 0.492
B-SZZ  (Falleri et al., 2014) 901 0.373 0.550 0.445
V-SZZ  (Bao et al., 2022) 1169 0.782 0.889 0.832
NVD  (NVD, 2009) 224 0.699 0.424 0.528
V0Finder  (Woo et al., 2021) 1374 0.834 0.745 0.787
Vercation Submission 1612 0.892 0.958 0.924
5.1.2. Comparison with SZZ algorithms
The basic idea of SZZ algorithms is to backtrace the commit history to locate the earliest commit that introduces the deletion lines removed by security patches. Regarding precision, recall, and F1-score, Vercation improves the best-performing baseline V-SZZ by 11.0%, 6.9% and 9.2%. And the difference between the F1-score of B-SZZ and AG-SZZ is small (0.413 vs. 0.492). We affirmed that there are two main causes of false alarms in this approach:
i) The line mapping algorithm based on edit distance does not consider the semantics of the source code. As shown in Listing LABEL:list:false_reason2, The deleted lines were encapsulated within the added functions, and in fact, they have the same semantics. Line mapping algorithms failed in these cases, unable to identify the true vic.
ii) The SZZ algorithm treats each deleted line as an independent origin for backtracking, resulting in multiple vic being identified. For example, the security patch of CVE-2020-35965 (cve, 2021) from FFmpeg has 5 deleted lines. The true vulnerability logic lies in the lack of checking the size of ymax before executing the memset zero operation, potentially resulting in out-of-bounds writes to memory. Our method utilizes a weighted allocation approach to increase the weight of memset operation, thus backtracing to the correct vic. However, SZZ algorithms separately backtrace different deleted lines, resulting in the identification of 5 different vic, significantly reducing the precision.
5.1.3. Comparison with V0Finder
In the same dataset, V0Finder achieved a precision of 83.4% and a recall of 74.5%. There are two reasons for the inaccuracies of V0Finder: 1) V0Finder generates a hash value for the entire vulnerable function as a vulnerability fingerprint. This fingerprint is coarse because the vulnerable function contains a lot of vulnerability-irrelevant code, introducing a significant amount of noise. 2) V0Finder considers the function containing all deleted patch code as a vulnerable function clone. This approach also overlooked the inserted code of security patch.
Accuracy of Vercation. Though Vercation performed well in the version identification, it still failed to identify inducing commits for a small number of vulnerabilities. One reason is that in some cases, slight modifications in the AST have little impact on the similarity between the AST before and after the modification, but are crucial in introducing vulnerabilities. Another reason is that despite using the weight allocation strategies, there are still some statements related to patch variables that are not relevant to the vulnerability but have been assigned high weights. The above two reasons may incorrectly identify vic, leading to a misidentification of the vulnerable version range.
RQ1: Vercation demonstrates superior performance in vulnerable version identification on our ground-truth dataset, achieving an F1 score of 92.4%. This significantly outperforms existing methods, including NVD reports, SZZ algorithms, and V0Finder. Compared to the best-performing baseline V-SZZ, Vercation shows a 9.2% improvement. These results highlight Vercation’s effectiveness in addressing limitations of current approaches, providing a more accurate tool for identifying vulnerable versions in open-source software.
5.2. RQ2: Ablation Study
We conduct ablation experiments to investigate the effectiveness of being harnessed by LLMs in terms of vulnerability comprehension task and compare the performance with only using the static analysis tool Joern parser.
Table 4. Performance of LLMs in Vulnerability Logic Comprehension.
Tool/Model Prompt Strategy Precision Recall F1 score
Joern - 0.426 0.574 0.489
Joern+GPT-4 Zero-shot 0.708 0.813 0.757
Joern+GPT-4 Few-shot 0.875 0.926 0.899
Joern+GPT-4 Few-shot+CoT 0.892 0.958 0.924
Joern+CodeLlama Zero-shot 0.683 0.780 0.728
Joern+CodeLlama Few-shot 0.863 0.911 0.886
Joern+CodeLlama Few-shot+CoT 0.868 0.938 0.901
5.2.1. Compared with Only Using Joern parser
As depicted in Table  4, the effectiveness of Joern in extracting vulnerability logic is significantly inferior to hardnessed by the power of LLM models, with an F1 score 45.4% less than Joern+GPT-4 (Few-shot+CoT). This discrepancy primarily stems from Joern’s reliance on fixed dependency extraction patterns, leading to the retrieval of numerous statements irrelevant to vulnerabilities. Meanwhile, we found that Joern cannot accurately address pointer structure dataflows, thus missing some statements that have data or control dependencies. To elaborate, on average, Joern extracts 3.12 and 2.71 times more vulnerable statements per CVE compared to combined GPT-4 and Codellama, respectively.
5.2.2. Performance of different LLM models
Table  4 illustrates the performance of combining the dangerous flows extracted by the Joern parser with GPT-4 and CodeLlama-13B models on our dataset, with the ”Prompt” column delineating different prompting strategies. Our observation reveals that GPT-4 prompted with Few-shot and CoT surpasses other models. Specifically, GPT-4 (Few-shot+CoT) exhibits 2.3% higher F1 score compared to the CodeLlama model. Moreover, the performance trends with GPT-4 remain consistently upward across different prompting strategies. When prompted with the Few-shot prompt, GPT-4 achieves a 14.2% higher F1 score than the Zero-shot prompt. This is further improved by 2.5% F1 with the addition of CoT prompting. We observed utilization of the Zero-shot prompt often results in an increased generation of false positive vulnerable statements, including the statements predefining patch variables (e.g., Line 9 in Listing 1). Conversely, in examples prompted by Few-shot, we prioritize understanding the triggering mechanisms of vulnerabilities, thereby facilitating LLM in learning a more accurate method of generating vulnerability logic.
RQ2: Harnessed by the power of LLMs, Vercation can generate the vulnerability logic and extract vulnerable codes from dangerous flows more precisely than traditional static analysis tools. The prompt strategy combining Few-shot and CoT techniques performs the best.
5.3. RQ3: Accuracy of Code Clone Detection
Here we evaluate the accuracy of the semantic-level similarity comparison technique utilized in P2.
5.3.1. Threshold Sensitivity
We used $\vartheta_{1}$ and $\vartheta_{2}$ to represent the threshold for syntactic similarity and semantic similarity respectively. $\vartheta_{3}$ represents the similarity_score used for the backtrace termination condition. To measure the sensitivity of the thresholds, we incrementally increased $\vartheta_{1}$, $\vartheta_{2}$ and $\vartheta_{3}$ by 0.1 from 0 to 1, and evaluated the identification results of each CVE in ground-truth. In addition, we evaluated each threshold with the metric F1-score, as we focused on both precision and recall. As shown in the results of Figure 4, the highest accuracy is achieved when $\vartheta_{1}$ is 0.9, $\vartheta_{2}$ is 0.8 and $\vartheta_{3}$ is 0.7. Specifically, when the thresholds are greater than the optimal values, the precision is higher while the recall decreases. In contrast, the precision slightly decreases.
Figure 4. Threshold Sensitivity of Vercation.
5.3.2. Accuracy of Semantic-level Code Clone Detection
Current methods primarily focus on function and file-level analysis[6,7], using coarse-grained representations like identifier sequences and control-flow graphs. While Vercation operates at sentence level, providing more nuanced analysis, differs significantly from existing code clone detection techniques. These conventional methods are not well-suited for sentence-level scenarios. Therefore, we compare some typical sentence-level code similarity detection methods with the method in Vercation’s P2. Table 5 summarizes the accuracy of different code clone detection methods.
1) Edit distance comparison calculates the minimum number of edit operations (insertions, deletions, substitutions) required to transform one code sentence into another. A smaller edit distance indicates a higher similarity between two code sentences.
2) Hash values comparison generates hash values for each code sentence and then compares the similarity of these hash values. Identical or similar hash values suggest that the code sentences may be similar.
3) CodeBERT embedding comparison uses the CodeBERT to generate embedding representations for each code sentence, then compute the cosine similarity between these embeddings. A cosine similarity closer to 1 indicates a higher semantic similarity between the code sentences.
We set the syntax similarity threshold $\vartheta_{1}$ for Edit distance and Hash values comparison, and the semantic similarity threshold $\vartheta_{2}$ for CodeBERT embedding comparison. The results show that the AST-based code clone detection perform the best performance. Because the two syntax-based statement clone detection methods terminate the backtrace in cases of low syntactic similarity without considering semantic similarity, thus leads to a large number of false negatives. The results show that our method is robust against low syntactic similarity. CodeBERT embedding comparison, while effective for semantic similarity, cannot capture fine-grained structural information. AST representations preserve the exact syntactic structure of code, allowing for more precise comparisons of code organization and logic flow. This structural analysis can identify clones that CodeBERT might miss, especially in cases where similar functionality is implemented with different coding patterns or variable names, which can be addressed by AST normalization.
Table 5. Accuracy of Code Clone Detection Methods.
Methods Level Precison Recall
Edit distance comparison syntax 0.85 0.84
Hash values comparison syntax 0.83 0.80
CodeBERT embedding comparison semantic 0.84 0.87
AST-based comparison syntax & semantic 0.89 0.96
5.3.3. Analysis of Refactoring Commits
In the backtrace of our dataset, we conducted a total of 276 comparisons between pre-commit and post-commit. On average, it takes 3.73 comparisons to identify the vic for each pc. During the commit backtrace, cases with significant syntactic differences ($\theta_{1}<0.9$) but minor semantic differences ($\theta_{2}\geq 0.8$) were classified as refactoring commits. Our analysis indicates that out of the 276 commit comparisons, 89 commits (32.2%) involved code refactoring. Furthermore, our analysis reveals that cases without any intermediate commit between the vic and the pc are comparatively uncommon, constituting only 11% of instances in our dataset.
RQ3: Refactoring commits occupy a significant proportion of the code changes history. The AST-based similarity comparison in Vercation performs better in the task of code changes detection than the singular syntactic and singular semantic analysis, significantly improves the precision of vulnerability-introducing commit identification, enhancing the overall effectiveness of Vercation in vulnerable version detection.
5.4. RQ4: Application of Vercation
We created a new vulnerability dataset consisting of 342 CVEs from the NVD. By developing a simple crawler, we obtained the corresponding CPE and CVE patches. We applied Vercation to the vulnerability dataset. The results indicate that the CPE for 134 CVEs is incomplete. Among them, 108 CVEs only reported a single vulnerable version, significantly increasing the risk of vulnerability propagation. Furthermore, the CPE for 56 CVEs is affected by the overinclusion problem. This is because NVD considers all versions before the OSS version mentioned in the vulnerability report as susceptible to attack, without conducting a detailed analysis. Even worse, we have discovered 4 CVEs that do not provide information about the vulnerable version.
(a) Box Plot for Introduce-To-Fix Delay of Top10 OSS.
(b) Box Plot for Fix-To-Publish Delay of Top10 OSS.
Figure 5. Time Delay in Different Periods of OSS Vulnerabilities.
Vulnerability Lifecycle Analysis. We selected the top 10 OSS in the NVD dataset based on their popularity on GitHub (the count of stars) and conducted the delay statistics at different periods. Figure 5(a) presents a box-plot of the introduce-to-fix time delay for the vulnerabilities. Due to the fact that code auditors often do not promptly conduct vulnerability risk analysis on newly added code, the introduce-to-fix time is usually quite lengthy, with some cases even lasting up to ten years. We found that Binutils has the lowest median delay (143 days), while Libtiff has the highest median delay (2132.5 days). As shown in Figure 5(b), the fix-to-publish time delay is a relatively short period. The median delay of FFmpeg, OpenJPEG, OpenVPN, Curl, php is within 30 days. However, Binutils and Libarchive have the highest median delay (158 and 149 days), so we urgently need to remind the developers of Binutils and Libarchive to fix the discovered vulnerabilities promptly. In these vulnerabilities, the fixes were often committed before the CVEs were published, as the vulnerability discoverers usually submit them to the OSS developers first. However, we found in 4 cases, the CVE was published before a fix had been committed, which greatly increases the risk of the vulnerability being exploited. This analysis highlights a critical window of vulnerability exposure in real-world projects. By accurately identifying the vic, Vercation enables developers to pinpoint when vulnerabilities were introduced, potentially speeding up the process of vulnerability identification and resolution.
Vulnerability Severity. We collected all the CVSS scores of NVD dataset. CVEs with CVSS Score in 4.0-6.9 (Medium) and 7.0-8.9 (High) accounted for the most, 49.7% and 49.1%. There are 2 CVEs that are critical (CVSS Score in 9.0-10.0) vulnerabilities that cause serious security issues, which are distributed in FFmpeg and php. The result indicates that the impact of OSS vulnerabilities is severe, we need to pay more attention to OSS vulnerabilities.
RQ4: Applying Vercation to real-world CVEs revealed significant practical impact. It identified 134 CVEs (38.1%) with incomplete or inaccurate CPE information in NVD reports. Analysis of the top 10 OSS projects showed substantial delays in vulnerability resolution, with some cases lasting up to ten years. This demonstrates Vercation’s potential to enhance vulnerability management in real-world scenarios by providing more accurate version information and insights into vulnerability lifecycles.
5.5. Threats to Validity
Limitations of Joern Parser. Vercation utilize Joern parser to generate the code property graph, containing abstract syntax tree, control flow graph and data dependency graph. However, we encountered some challenges with certain C/C++ constructs that Joern did not handle well, impeding its ability to track data flow and missing essential constructors correctly. Additionally, Joern may face difficulties in outputting data dependencies related to pointers, particularly in languages like C/C++ where pointers are heavily used. This limitation could potentially affect the efficacy of our approach.
Generalizability of the Dataset. Our study primarily focused on C/C++ open-source software projects. While these are widely used languages, the effectiveness of Vercation on other programming languages or closed-source software remains untested. The vulnerability patterns and code structures in other languages might differ, potentially affecting the performance of our approach. Future work could involve extending Vercation to support a broader range of programming languages and software types.
LLM Variability and Reproducibility. As LLMs continue to evolve rapidly, the performance of Vercation might change with newer versions of these models. This could affect the long-term consistency of our results. Additionally, LLMs may have biases or limitations in their training data that could influence their ability to understand certain types of vulnerabilities or code structures. To mitigate this, we have provided detailed information about the LLM versions and prompts used in our study, but future research may need to account for potential variations in LLM capabilities and performance as these models continue to develop.
6. Discussion
Time Efficiency. In our experimental setup, we took 10 days to build the software and version pool. Then we took 30 days to build the ground-truth dataset, including the time taken to establish the dependent environment for each software version, as well as the time to perform vulnerability verifying and labeling. Vercation analyzes each vulnerability on an average of 28.61 seconds.
Branch Fixing Oversights. In some cases, we observe that security fixes are promptly incorporated into the latest version branches of OSS. However, rectifying oversights frequently transpires during the backporting of fixes to older release branches. In such scenarios, projects not utilizing the latest branch version may encounter a deficiency in receiving security updates. Simultaneously, the latest versions updated in older branches remain exposed to risks.
Limitation and Future Work. Vercation operates under certain assumptions that may limit its applicability. Firstly, it is tailored for C/C++ projects, yet its methodology can be applied to other programming languages. With an accurate exploration of common vulnerability types in diverse languages, Vercation could identify vulnerable versions in those languages. Secondly, Vercation exclusively addresses vulnerabilities within functions, thus omitting inter-functional vulnerabilities. However, we are contemplating the integration inter-functional data-flow information into the vulnerability signatures to tackle such scenarios. In the evaluation results presented in section 5.2, Vercation demonstrated notable precision (93%) and recall (98%). There exists room for further enhancement in both precision and recall. In semantic similarity comparisons, the utilization of heavyweight program analysis techniques, such as symbolic execution, could enhance the understanding of code semantics, thereby reducing false negatives.
7. Related work
Vulnerability Fixing Commit Analysis. Several studies have focused on analyzing vulnerability fixing commit (VFC) to extract vulnerability features, which can be used for various purposes such as vulnerability detection, classification, and affected version identification. VulPecker (Li et al., 2016) extracts code patterns from VFC to detect similar vulnerabilities in other software versions. SySeVR (Li et al., 2021) uses a systematic approach to extract syntax- and semantic-based vulnerability features from patches. VulDeePecker (Li et al., 2018) leverages deep learning techniques to learn vulnerability patterns from code gadgets extracted from vulnerability fixing commits. VulCNN (Wu et al., 2022) uses convolutional neural networks to learn vulnerability patterns from code changes in VFC. Most existing VFC analysis methods rely heavily on predefined rules or patterns, which may not capture the full semantic context of vulnerabilities. They also often struggle with handling complex code changes and may miss subtle vulnerability indicators. Vercation incorporates advanced VFC analysis by leveraging LLMs to understand the semantic context of code changes in fixing commits, potentially capturing more nuanced vulnerability features.
Vulnerable Detection for Open-source Software. Several approaches have been proposed for vulnerable version identification. V-SZZ (Bao et al., 2022) uses the SZZ algorithm to backtrace patch commits and identify vulnerability-introducing changes. MVP (Xiao et al., 2020) employs program slicing to extract vulnerability and patch signatures, then uses these to identify potentially vulnerable functions. AFV (Shi et al., 2022) introduces a vulnerability-centric approach for precise analysis of affected versions in web vulnerabilities.V0Finder (Woo et al., 2021) generates fingerprints for vulnerable functions and uses a clone-based technique to detect vulnerable versions across different software releases. VCCFinder (Perl et al., 2015) uses machine learning techniques to identify vulnerability-contributing commits. VUDDY (Kim et al., 2017) proposes a scalable approach for vulnerable code clone detection in large-scale code bases. Existing approaches often rely on predefined patterns and syntactic-level analysis, leading to imprecision in vulnerability characterization and code clone detection. Vercation combines static analysis with LLM for semantic understanding and introduces AST-based semantic-level code clone detection.
Code Refactoring Detection. Some researchers conducted studies on the characteristics of code refactoring not altering the semantic behavior of the code and have proposed detection methods. MLRefScanner (Noei et al., 2024) detects refactoring commits in machine-learning Python projects by analyzing commit history and extracting features that represent refactoring activities without altering semantic behavior. Abid et al. (Abid et al., 2020) perform a study outlining different levels of refactoring from code-level to architecture and discuss the importance of maintaining behavior during refactoring. Eman et al. (AlOmar et al., 2021) explores refactoring practices in industrial settings, focusing on how refactoring is integrated into the code review process while ensuring that the original software behavior is preserved. Existing code refactoring detection methods primarily focus on high-level changes like renaming, method/class moves. Our approach, however, targets sentence-level changes through fine-grained AST expansion and normalization, including inline function expansion during AST generation. This level of detail in internal code structure analysis is a novel aspect not previously considered.
8. Conclusion
This paper proposed Vercation, an approach designed for vulnerable version identification of open-source C/C++ software. The core idea involves a symbiotic combination of static analysis, LLM, and semantic-level code clone detection. Our approach introduces two key innovations: leveraging LLMs to enhance extraction of vulnerability-related statements and employing semantic-level code clone detection based on expanded Abstract Syntax Trees (ASTs). Experimental results on our dataset of 74 CVEs across 11 popular open-source projects validate that Vercation surpasses existing techniques for vulnerable version identification, demonstrating notable improvements in both precision and recall. Equipped with the discovery from Vercation, developers can effectively address potential threats arising from cross-version vulnerabilities in OSS, thereby enhancing the overall security of the software ecosystem.